What is the volume of a 0.19 M sulfuric acid (H2SO4) solution that contains 0.34 mol sulfuric acid?
1
Expert's answer
2020-03-09T09:40:52-0400

The molarity can be defined as following:

M = n / V

where M - molarity of the solution, n - number of moles of the solute, V - volume of the solution.

From here:

V = n / M = 0.34 mol / 0.19 M = 1.79 L


Answer: 1.79 L